OpenStudy (anonymous):

Solve cos2θ + cosθ = 0 for 0° ≤ θ < 90°.

OpenStudy (anonymous):

Use Cos 2 theta = 2 Cos^2 theta -1 to get a quadratic in theta.

OpenStudy (anonymous):

COS2A = 2COSA^2 - 1 PUT THAT INTO THE EQUATION 2COSA^2 +COSA -1 =0 PUT COSA = X 2X^2 +X -1 =0 2X(X +1) -(X+1)=0 X= -1 , I/2 COSA = -1 , 1/2 A= pi, pi/3

OpenStudy (anonymous):

COSA^2 IS FOR (COS^2A)